# How connection probability shapes fluctuations of neural population dynamics

### v1

@@ IV.5.1 Dependence on the connection probability
in the interval $(0,1)$ if $-2/3<\frac{2\tau C}{-w\mu_{0}}\left(\beta^{-2}-\sigma_{\text{ext}}^{2}/4\right)<1/2$. In the limit of a steep single-neuron transfer function ($\beta\to\infty$) and in the absence of external noise, the maximum is attained at $p_{\text{max}}=2/3$. Such maximum is indeed verified by simulations (Fig. 7b, right). Furthermore, the variance $\text{var}(\bar{h})$ approaches zero for $p\to 0$ and approaches a non-vanishing value for $p\to 1$.
In contrast to the 2nd-order theory, the 1st-order MF model cannot explain the strictly convex and non-monotonic behavior of the variances of the population rates and input potentials, respectively (Fig. 7b). In the second scenario with constant $C$, the 1st-order MF theory is obtained from the above formulas by fixing the ratio $p/C$ and letting $p\to 1$. Thus, the common factor $\bigl{(}\frac{\hat{\mu}_{0}p}{\hat{\tau}C}+\hat{\sigma}_{\text{ext}}^{2}\bigr{)}$ in Eqs. (46) and (48) remains unchanged, whereas the second factor becomes independent of $p$. Because the common factor is an increasing linear function of $p$, the variances are thus also increasing linear functions, in line with simulations of the 1st-order mean-field model (but in contradiction to microscopic simulations and 2nd-order MF theory).
@@ IV.5.2 Dependence on the mean stimulus strength
We started our paper with network simulations pointing out marked quantitative and qualitative discrepancies of the mean-connectivity network with respect to the stimulus dependence of the rate variability (Fig. 1). In particular, we observed a strong overestimation of the rate variance by the mean-connectivity network compared to the original quenched network. Furthermore, for a noisy external stimulus and large network size ($N=50000$), we observed a suppression of the rate variability by an increase in the mean stimulus strength in the quenched network. In contrast, the mean-connectivity network incorrectly predicted an increase in rate variability rather than a suppression. As we will see now, these observations can be well explained by our 2nd-order MF theory.
To this end, we study two opposing scenarios: first, we consider a small (“mesoscopic”) network without noise in the external stimulus ($\sigma_{\text{ext}}=0$). In this case, the rate variability solely originates intrinsically from the finite-size fluctuations. Second, we consider a large (“macroscopic”) network with a noisy external stimulus ($\sigma_{\text{ext}}>0$). In this case, the rate variability is purely externally generated. In both cases, we find that the first-order MF theory largely overestimates the variance of the population rate in the quenched and annealed network, whereas the 2nd-order MF theory, Eq. (43), correctly predicts the magnitude of the rate variance (Fig. 8a,b). Because the 1st-order MF model is the exact mean-field model for the mean-connectivity network, this large deviation explains our observation in Fig. 1.
@@
These observations can be understood analytically using the approximation Eq. (46), (47). Let us first discuss how the factor $\hat{F}_{h}$ in Eq. (46) depends on $\hat{\mu}_{0}$. For $p=1$, i.e. in the 1st-order MF theory, $\hat{F}_{h}$ is exactly given by the function $Q(\hat{\mu}_{0})$ displayed by the dotted line in Fig. 8c. This function, and thus $\hat{F}_{h}$, grows monotonotically for $\hat{\mu}_{0}>0$ until it reaches a maximum at $\hat{\mu}_{0}^{*}=-\mu_{0}^{*}/(wr_{\text{m}})=0.5$, which corresponds to the upper bound $r_{\text{m}}/2$ of the biologically relevant dynamical range. For $p<1$, the factor $\hat{F}_{h}$ is the quotient of the symmetric function $Q$ and an increasing function, which shifts the maximum of $\hat{F}_{h}$ to a value $<1/2$ (Fig. 8c, solid line). Hence, $\hat{F}_{h}$ exhibits a non-monotonic behavior in the biologically relevant range $0<\mu_{0}<\mu_{0}^{*}$. Besides the shift of the maximum to smaller values of the mean stimulus strength $\hat{\mu}_{0}$, the maximum also decreases when the connection probability is lowered from $p=1$ to smaller values. This decrease is a consequence of dividing the function $Q(\hat{\mu}_{0})$ by the square-root term that is strictly larger than one for $p<1$. Importantly, $\hat{F}_{h}$ can decrease by a large factor if the single-neuron transfer function is steep, i.e. $\hat{\beta}\gg 1$, which explains the drastic decrease of the rate variance from 1st- to 2nd-order MF theory (Fig. 8a,b, where $\hat{\beta}=500$).
@@
The right-hand side is a product of a monotonically increasing function of $\hat{\mu}_{0}$ and the function $Q(\hat{\mu}_{0})$, which is maximized at $\hat{\mu}_{0}^{*}=1/2$. Therefore, the maximum of $\hat{\mu}_{0}\hat{F}_{h}$ is shifted to values $\hat{\mu}_{0}>1/2$ for all $p\in[0,1]$ (Fig. 8d). As a consequence, the variance of the population rate $\text{var}(r)$ is a monotonically increasing function of the mean stimulus strength in the biologically relevant range $20\leq\hat{\mu}_{0}\leq 1/$ 2 (Fig. 8a right, $\mu_{0}<50$ mV), and thus an increase of the mean stimulus strength always increases the rate variability in the absence of external common noise (Fig. 8a left).
To keep the arguments simple, we studied here two clear-cut cases – one with and another without external noise. According to our theory, Eq. (45), the rate variance is a sum of two terms, where each term on its own corresponds to one of these cases. In general, the theory also applies to the mixed case, where the rate variance is the sum of externally and intrinsically induced variability. For example, in the simulations of Fig. 1, we also compared two networks of size $N=1000$ and $N=50000$ but the external common noise was present in both cases. Nonetheless, the mesoscopic case of $N=1000$ with external noise can still be qualitatively understood by the limit case $\sigma_{\text{ext}}=0$ because the finite-size fluctuations dominate over the external noise.
@@ V Conclusions
In this work, we set out to derive a low-dimensional, stochastic population dynamics for random recurrent networks of spiking neurons that correctly accounts for fluctuations when the network is neither fully connected ($p=1$) nor infinitely large ($N\to\infty$). While such regime is biologically highly relevant, to our knowledge it has not yet been systematically investigated within the frameworks of previously developed mean-field theories [29, 51]. By deriving a low-dimensional, stochastic mean-field model that allows one to study specifically the combined population effects of quenched disorder and stochastic neuronal firing in spiking networks, we thus hope to close an important gap in the use cases of simplified descriptions of networks and to provide the foundations for a mean-field modeling framework of neural variability at the mesoscopic population level.
@@
The analysis of the new “second-order” mean-field model showed that the disorder of the connectivity for $p<1$ has a drastic effect on the neural variability, the response to time-dependent stimuli, and the stability of the network in biologically relevant regimes. Specifically, our main findings with regard to the effects of a finite connectivity $10<p<$ 1 and finite network size $N<\infty$ are the following: (i) A finite variance in the input currents ($\sigma^{2}>0$) caused by the disorder of a non-full, random connectivity with finite $C$ leads to an effective broadening of the neuron transfer function, but the stationary mean firing rate remains rather unaffected (Fig. 3). (ii) In the presence of synaptic delays, the stability properties of fixed points of the dynamics and locations of bifurcations can change considerably. In particular, we showed that the connectivity disorder can stabilize networks in regimes where a “first-order” mean-field theory corresponding to a mean-connectivity approximation would predict oscillatory dynamics (Fig. 4). (iii) The population-rate response to high-frequency stimuli and (iv) the variance of the population rate are significantly lower than predicted by a “first-order” mean-field theory that does not correctly account for the incoherent fluctuations caused by the connectivity disorder (Fig. 5 and 7). Moreover, (v) in the presence of shared noise due to common external input, an increase in the mean external input can actually lead to a decrease of the variance of the population rate (Figs. 1 and 8) – an effect which again is not captured in a “first-order” mean-field theory. Overall, we provided simple analytical explanations for the effects of finite connectivity and finite network size in terms of the slope of the population transfer function and the multiplicative character of the finite-size noise. In particular, we found that a decreasing connection probability $p$ lowers the slope $F_{h}$ through an increase of the membrane potential variance $\sigma^{2}$ and that, at finite network size $N$, the neural variability has a rate-dependent contribution that is proportional to the product of mean stimulus strength (or firing rate) and the slope of the transfer function (Fig. 8d, Eq. (45)).
@@
In our “dynamically annealed” approximation, we assumed that each presynaptic spike would be randomly distributed among all neurons according to the connection probability $p$, independent of the actual, fixed (quenched) synaptic connectivity matrix of a given network. A systematic comparison of our theory with simulations of quenched networks showed that our second-order mean-field theory captures reliably the behavior of quenched networks with respect to the main findings described above. Deviations of the annealed network (from the quenched network) and, correspondingly, the 2nd-order MF model, are noticeable for several statistics, though. However, these deviations were considerably smaller than the respective deviations of the mean-connectivity network and the corresponding 1st-order MF model used in previous studies [49, 35, 50]. The remaining deviations of the annealed network can most likely be attributed to the neglect of temporal correlations in the recurrent input caused by the annealing, i.e. the incessant resampling of the connections in time. This explanation would be consistent with the underestimation of the low-frequency power by the annealed network which we observed in Fig. 6. For infinitely large, sparsely connected networks, a proper account of temporal correlations can, in principle, be obtained through a self-consistent treatment of the auto-correlations of the recurrent fluctuations [68, 56], which typically adds low-frequency power [56]. Such self-consistent theory would lead to colored noise rather than white noise for the incoherent fluctuations in Eq. (12). However, how to build a stochastic, dynamical population model for finite-size, random networks that accounts for temporal correlations in a self-consistent manner is an open theoretical problem which goes beyond the scope of the present study.
@@
Interestingly, perhaps not surprisingly, the comparison of our theory with simulations of “dynamically annealed” networks shows that it can be considered almost exact for the annealed case. One may ask whether a biological realization of such a “dynamically annealed” network, where our theory becomes exact, might actually exist. A possible origin of such randomly distributed spikes might be the probabilistic vesicle release: A fully connected network where synapses would transmit a spike with probability $p$ would be an exact realization of the system we describe by our approximation. Thus, the annealed network model may also be regarded as a simple caricature of a spiking neural network with probabilistic synaptic transmission. Another type of annealing occurs in networks with full connectivity but dynamic synapses (short-term synaptic plasticity). Recently, such dynamically changing networks have also been reduced to low-dimenensional Langevin dynamics based on a similar Poisson neuron model [69, 70].
@@
For the sake of a systematic and thorough development of the theory, we here began with the case of one population corresponding to a single neuron type; especially, we focused on the case of an inhibitory network. However, our theory suggests a straightforward extension to multiple populations. Cortical networks typically consist of several cell types and are generally described as excitatory-inhibitory networks that are, in the simplest case, organized as two coupled populations. Rate models of such two-population E-I networks remain a popular tool to describe network dynamics up to this day, whether to characterize inhibition stabilization in cortical networks [71] or spatial patterns of oscillatory activity observed in motor cortex [50, 72]. It will certainly be worthwhile to investigate how the effects of finite connectivity and network size we began to explore here may shape the more complicated dynamics of networks that comprise more than one population. Furthermore, a multi-population extension would offer the possibility to model specific cortical microcircuits, such as canonical circuit models [73] or cortical column models with multiple interneuron types [43], on the mesoscopic scale, which accounts for fluctuations.
@@
For simulations and concreteness, we made some further specific modeling choices: For example, we studied random networks with fixed in-degree, i.e. an identical number of presynaptic connections, and chose a specific transfer function (or $f-I$ curve) $\phi(h)$ that relates the instantaneous firing rate to the intensity or input current $h$. While our specific choice of the transfer function also enabled us to analytically evaluate the effect of a finite variance in the input currents on the population rate, the reduction of the spiking network dynamics to a low-dimensional system of coupled SDEs should remain valid for arbitrary choices of the transfer function, albeit probably more difficult to compute. The assumption of fixed in-degree allowed us to assume that the mean input current as well as the strengths of the fluctuating parts are identical among neurons, but in practice we observed that a purely random Erdös-Réniy network shows almost identical statistics as those described in the previous sections for networks with fixed in-degree.
Here, we showed how to derive low-dimensional stochastic models that describe fluctuations in a complex, biological model system, namely a finite-size spiking neural networks with non-full connectivity. Already in the simple case of one population and Poissonian spiking, we found highly non-trivial effects such as the non-monotonic dependence of the variance of the recurrent input on connection probability (Fig. 7b) and the weak suppression of variability by external stimuli (Fig. 8b). Thus, even though a single population and Poisson spiking dynamics may be too simple to quantitatively model certain phenomena in cortex (such as a stronger suppression of variability [39]), the fact that our 2nd-order MF theory explains non-trivial fluctuation effects demonstrates the power of the approach as a proof of principle. Importantly, with the biological extensions discussed above, we believe that our theory may pave the way for a useful modeling framework for variability in cortical circuits at the mesoscopic population level.

### v2

@@ IV.5.1 Dependence on the connection probability
in the interval $(0,1)$ if $2/3<\frac{2\tau C}{-w\mu_{0}}\left(\beta^{-2}-\sigma_{\text{ext}}^{2}/4\right)<1/2$. In the limit of a steep single-neuron transfer function ($\beta\to\infty$) and in the absence of external noise, the maximum is attained at $p_{\text{max}}=2/3$. Such maximum is indeed verified by simulations (Fig. 5b, right). Furthermore, the variance $\text{var}(\bar{h})$ approaches zero for $p\to 0$ and approaches a non-vanishing value for $p\to 1$.
In contrast to the 2nd-order theory, the 1st-order MF theory cannot explain the strictly convex and non-monotonic behavior of the variances of the population rates and input potentials, respectively (Fig. 5b). In the second scenario with constant $C$, the 1st-order MF theory is obtained from the above formulas by fixing the ratio $p/C$ and letting $p\to 1$. Thus, the common factor $\bigl(\frac{\hat{\mu}_{0}p}{\hat{\tau}C}+\hat{\sigma}_{\text{ext}}^{2}\bigr)$ in Eqs. (47) and (49) remains unchanged, whereas the second factor becomes independent of $p$. Because the common factor is an increasing linear function of $p$, the variances are thus also increasing linear functions, in line with simulations of the 1st-order mean-field model (but in contradiction to microscopic simulations and 2nd-order MF theory).
@@ IV.5.2 Dependence on the mean stimulus strength
We started our paper with network simulations pointing out marked quantitative and qualitative discrepancies of the mean-connectivity network with respect to the stimulus dependence of the rate variability (Fig. 1). In particular, we observed a strong overestimation of the rate variance by the mean-connectivity network compared to the original quenched network. At first sight, it may appear counter-intuitive that more homogeneous connectivity leads to a stronger variance in the population rate. This apparent paradox results from the fact that in mean-connectivity networks, finite-size fluctuations are ‘directly’ picked up by the steep single-unit transfer function that also determines the population rate, $F(h,\sigma^{2})=\phi(h)$ for $\sigma^{2}=0$, whereas the effect of these fluctuations at the population level is attenuated for networks with heterogeneous connectivity due to due to the flattening of the effective population rate transfer function; the difference in the population rate variance can then be seen from Eq. (46). We will make this argument more explicit below. We also note that the large variance of the population rate in the mean-connectivity network (or MF1) can be attributed to an increased high-frequency power (Fig. 4b).
Furthermore, for a noisy external stimulus and large network size ($N=50000$), we observed a suppression of the rate variability by an increase in the mean stimulus strength in the quenched network. In contrast, the mean-connectivity network incorrectly predicted an increase in rate variability rather than a suppression. As we will see now, these observations can also be well explained by our 2nd-order MF theory.
Let us study two opposing scenarios: first, we consider a small (“mesoscopic”) network without noise in the external stimulus ($\sigma_{\text{ext}}=0$). In this case, the rate variability solely originates intrinsically from the finite-size fluctuations. Second, we consider a large (“macroscopic”) network with a noisy external stimulus ($\sigma_{\text{ext}}>0$). In this case, the rate variability is purely externally generated. In both cases, we find that the first-order MF theory largely overestimates the variance of the population rate in the quenched and annealed network, whereas the 2nd-order MF theory, Eq. (43), correctly predicts the magnitude of the rate variance (Fig. 6a,b). Because the 1st-order MF dynamics is the exact MF dynamics for the mean-connectivity network, this large deviation explains our observation in Fig. 1.
@@
These observations can be understood analytically using the approximation Eq. (47), (48). Let us first discuss how the factor $\hat{F}_{h}$ in Eq. (47) depends on $\hat{\mu}_{0}$. For $p=1$, i.e. in the 1st-order MF theory, $\hat{F}_{h}$ is exactly given by the function $Q(\hat{\mu}_{0})$ displayed by the dotted line in Fig. 6c. This function, and thus $\hat{F}_{h}$, grows monotonically for $\hat{\mu}_{0}>0$ until it reaches a maximum at $\hat{\mu}_{0}^{*}=-\mu_{0}^{*}/(wr_{\text{m}})=0.5$, which corresponds to the upper bound $r_{\text{m}}/2$ of the biologically relevant dynamical range. For $p<1$, the factor $\hat{F}_{h}$ is the quotient of the symmetric function $Q$ and an increasing function, which shifts the maximum of $\hat{F}_{h}$ to a value <1/2<1/2 (Fig. 6c, solid line). Hence, $\hat{F}_{h}$ exhibits a non-monotonic behavior in the biologically relevant range $0<\mu_{0}<\mu_{0}^{*}$. Besides the shift of the maximum to smaller values of the mean stimulus strength $\hat{\mu}_{0}$, the maximum also decreases when the connection probability is lowered from $p=1$ to smaller values. This decrease is a consequence of dividing the function $Q(\hat{\mu}_{0})$ by the square-root term that is strictly larger than one for $p<1$. Importantly, $\hat{F}_{h}$ can decrease by a large factor if the single-neuron transfer function is steep, i.e. $\hat{\beta}\gg 1$, which explains the drastic decrease of the rate variance from 1st- to 2nd-order MF theory (Fig. 6a,b, where $\hat{\beta}=500$, corresponding to $\beta=5$mV-1 as in Fig. 1).
@@
The right-hand side is a product of a monotonically increasing function of $\hat{\mu}_{0}$ and the function $Q(\hat{\mu}_{0})$, which is maximized at $\hat{\mu}_{0}^{*}=1/2$. Therefore, the maximum of $\hat{\mu}_{0}\hat{F}_{h}$ is shifted to values $\hat{\mu}_{0}>1/2$ for all $p\in[0,1]$ (Fig. 6d). As a consequence, the variance of the population rate $\text{var}(r)$ is a monotonically increasing function of the mean stimulus strength in the biologically relevant range $0\leq\hat{\mu}_{0}\leq 1/2$ (Fig. 6a right, $\mu_{0}<50$ mV), and thus an increase of the mean stimulus strength always increases the rate variability in the absence of external common noise (Fig. 6a left).
To keep the arguments simple, we studied here two clear-cut cases – one with and another without external noise. According to our theory, Eq. (46), the rate variance is a sum of two terms, where each term on its own corresponds to one of these cases. In general, the theory also applies to the mixed case, where the rate variance is the sum of externally and intrinsically induced variability. For example, in the simulations of Fig. 1, we also compared two networks of size $N=1000$ and $N=50000$ but the external common noise was present in both cases. Nonetheless, the mesoscopic case of $N=1000$ with external noise can still be qualitatively understood by the limit case $\sigma_{\text{ext}}=0$ because the finite-size fluctuations dominate over the external noise. Lastly, we remark that as $\hat{\mu}_{0}\to 0$, e.g. for strongly inhibition-dominated networks, the population rate variance decreases for both the mean-connectivity network (MF1) and MF2 in both cases as can be seen from Fig. 6c,d.
@@ V Conclusions
In this work, we set out to derive a stochastic population dynamics for finite-size networks of spiking neurons that correctly accounts for fluctuations at the mesoscopic scale when the connectivity is random – as in the case of non-full connectivity ($p<1$). While such regime is biologically highly relevant, to our knowledge it has not yet been systematically investigated within the frameworks of previously developed mean-field theories [29, 30]. The low-dimensional, stochastic population model derived here allows one to study specifically how the combination of quenched disorder and stochastic neuronal firing affects the mesoscopic activity of finite-size neuronal populations. We thus hope to close an important gap in the use cases of simplified descriptions of networks and to provide the foundations for a mean-field modeling framework of neural variability at the mesoscopic population level.
@@
The analysis of the new “second-order” mean-field (MF2) dynamics showed that the quenched connectivity disorder for $p<1$ has a drastic effect on the neural variability, the response to time-dependent stimuli, and the stability of the network in biologically relevant regimes. Specifically, our main findings with regard to the effects of a finite connectivity $0<p<1$ and finite network size $N<\infty$ are the following: (i) A finite variance in the input currents ($\sigma^{2}>0$) caused by the quenched random connectivity with finite $C=pN$ leads to an effective broadening of the neuron transfer function, but the stationary mean firing rate remains rather unaffected (Figs. 3a, 8b). (ii) In the presence of synaptic delays, the stability properties of fixed points of the dynamics and locations of bifurcations can change considerably. In particular, we showed that the connectivity disorder can stabilize networks in regimes where a “first-order” mean-field (MF1) theory corresponding to a mean-connectivity approximation would predict oscillatory dynamics (Fig. 3b,c). (iii) The population-rate response to high-frequency stimuli and (iv) the variance of the population rate are significantly lower than predicted by the MF1 theory (Fig. 3d,e and 5). Moreover, (v) in the presence of shared noise due to common external input, an increase in the mean external input can actually lead to a decrease of the variance of the population rate (Figs. 1 and 6) – an effect which again is not captured in the MF1 theory.
Overall, we provided simple analytical explanations for the effects of finite connectivity and finite network size in terms of the slope of the population transfer function and the multiplicative character of the finite-size noise. In particular, we found that a decreasing connection probability $p$ lowers the slope $F_{h}$ through an increase of the membrane potential variance $\sigma^{2}$ and that, at finite network size $N$, the neural variability has a rate-dependent contribution that is proportional to the product of mean stimulus strength (or firing rate) and the slope of the transfer function (Fig. 6d, Eq. (46)).
Of note, our mesoscopic approach differs from previous theoretical studies that focussed on how quenched random connectivity shapes microscopic dynamics [48, 75, 72, 76]. Neural networks with quenched disorder can display a wide repertoire of dynamical states even under homogeneous random connectivity [76], as considered here. In particular, inhibition-dominated networks operating near the edge of chaos produce rich single-neuron dynamics that may support complex computations [77, 76]. Importantly, this critical dynamical state refers to the microscopic dynamics, as characterized by broad distributions of pairwise correlations [77, 76] and neuronal time scales [78]. The annealed approximation underlying our theory cannot reproduce these microscopic features. However, this limitation is not problematic for our mesoscopic description of the coarse-grained population activity. As already noted in [76], the mesoscopic population activity is largely unaffected by the critical dynamics at the microscopic scale. We have furthermore shown here that, at the mesoscopic scale, the annealed approximation captures surprisingly well both the dynamics and their fluctuations across the full range of connection probabilities.
@@
In our “dynamically annealed” approximation, we assumed that each presynaptic spike would be randomly distributed among all neurons according to the connection probability $p$, independent of the actual, fixed (quenched) synaptic connectivity matrix of a given network. A systematic comparison of our theory with simulations of quenched networks showed that our MF2 theory captures reliably the behavior of quenched networks with respect to the main findings described above. Deviations of the annealed network (from the quenched network) and, correspondingly, the MF2 dynamics, are noticeable for several statistics, though. However, these deviations were considerably smaller than the respective deviations of the mean-connectivity network and the corresponding MF1 dynamics used in previous studies [47, 36, 79].
The remaining deviations of the annealed network can most likely be attributed to the neglect of temporal correlations in the recurrent input caused by the annealing, i.e. the incessant resampling of the connections in time. This explanation would be consistent with the underestimation of the low-frequency power by the annealed network which we observed in Fig. 4. For infinitely large, sparsely connected networks, a proper account of temporal correlations can, in principle, be obtained through a self-consistent treatment of the auto-correlations of the recurrent fluctuations [75, 62], which typically adds low-frequency power [62]. Such self-consistent theory would lead to colored noise rather than white noise for the incoherent fluctuations in Eq. (12). However, how to build a stochastic, dynamical population model for finite-size, random networks that accounts for temporal correlations in a self-consistent manner is an open theoretical problem which goes beyond the scope of the present study.
Figure 7: Variance of $r$ and $\bar{h}$ for a network of LIF neurons with a rectified linear hazard function $\phi(h)=\beta_{\text{lin}}[h-\theta_{\text{lin}}]_{+}$ and a post-spike reset to $h_{\text{reset}}$ as a function of connection probability $p$. The LIF network behaves qualitatively similar to the Poisson network without reset in this paper (cf. Fig. 5). Variances are calculated from network simulations of quenched (green) and annealed (red) networks for (a) $N=1000$ and (b) $C=100$ for different connection probabilities $p$. The rectified linear hazard rate was matched both by value and slope of the single neuron transfer function Eq. (2) at a reference point $h_{\text{ref}}=-0.25$ mV. Parameters: $h_{\text{reset}}=-15$ mV, $\beta_{\text{lin}}\approx 91.32$ mV-1, $\theta_{\text{lin}}\approx-0.37$ mV, w=1w=-1 mVs, $\mu_{0}=10$ mV.
@@
Interestingly, perhaps not surprisingly, the comparison of our theory with simulations of “dynamically annealed” networks shows that it can be considered almost exact for the annealed case. One may ask whether a biological realization of such a “dynamically annealed” network, where our theory becomes exact, might actually exist. A possible origin of such randomly distributed spikes might be the probabilistic vesicle release: A fully connected network where synapses would transmit a spike with probability $p$ would be an exact realization of the system we describe by our approximation. Thus, the annealed network model may also be regarded as a simple caricature of a spiking neural network with probabilistic synaptic transmission. Another type of annealing occurs in networks with full connectivity but dynamic synapses (short-term synaptic plasticity). Recently, such dynamically changing networks have also been reduced to low-dimensional Langevin dynamics based on a similar Poisson neuron model [58, 59].
@@
For the sake of a systematic and thorough development of the theory, we here began with the case of one population corresponding to a single neuron type; especially, we focused on the case of an inhibitory network. However, our theory suggests a straightforward extension to multiple populations. Cortical networks typically consist of several cell types and are generally described as excitatory-inhibitory networks that are, in the simplest case, organized as two coupled populations. Rate models of such two-population E-I networks remain a popular tool to describe network dynamics up to this day, whether to characterize inhibition stabilization in cortical networks [80] or spatial patterns of oscillatory activity observed in motor cortex [79, 81]. It will certainly be worthwhile to investigate how the combined effects of random connectivity and network size we began to explore here may shape the more complicated dynamics of networks that comprise more than one population. Furthermore, a multi-population extension would offer the possibility to model specific cortical microcircuits, such as canonical circuit models [82] or cortical column models with multiple interneuron types [41], on the mesoscopic scale, which accounts for fluctuations.
@@
While a systematic analysis of the effect of resets is beyond the scope of this paper, we present here some first simulation results. We consider a quenched network, in which the input potential is reset to $h_{\text{reset}}$ after each spike. With reset, the single-neuron hazard rate no longer requires an upper bound, and we therefore adopt a rectified linear hazard rate [28]. As an example, we repeat the simulations from Fig. 5 with reset. The results (Fig. 7) show that the variances of the population rate $r$ and mean input potential $\bar{h}$ as a function of the connection probability $p$ behave qualitatively similarly to the case without reset (Fig. 5). Specifically, for fixed $N$, the variance of $r$ increases with $p$, while the variance of $\bar{h}$ decreases. For fixed $C$, the variance of $r$ increases, and the variance of $\bar{h}$ exhibits a maximum. Importantly, the annealed approximation now matches the quenched network much more closely than in the absence of reset. This aligns with our earlier explanation that discrepancies between the quenched network and the annealed approximation arise from neglecting temporal correlations in the incoherent recurrent input due to annealing. The reset reduces the enhanced low-frequency power of the model without reset, effectively whitening the spike-train power spectra and thus reducing temporal correlations in the input. In other words, there are less temporal correlations to be removed by the annealing, leading to a better fit.
For simulations and concreteness, we made some further specific modeling choices: For example, we studied random networks with fixed in-degree, i.e. an identical number of presynaptic connections, and chose a specific transfer function (or fIf-I curve) $\phi(h)$ that relates the instantaneous firing rate to the intensity or input current $h$. While our specific choice of the transfer function also enabled us to analytically evaluate the effect of a finite variance in the input currents on the population rate, the reduction of the spiking network dynamics to a low-dimensional system of coupled SDEs should remain valid for arbitrary choices of the transfer function, albeit probably more difficult to compute. The assumption of fixed in-degree allowed us to assume that the mean input current as well as the strengths of the fluctuating parts are identical among neurons, but in practice we observed that a purely random Erdős-Réniy network shows almost identical statistics as those described in the previous sections for networks with fixed in-degree.
Here, we showed how to derive low-dimensional stochastic models that describe fluctuations in a complex, biological model system, namely a finite-size spiking neural networks with random connectivity. Already in the simple case of one population and Poissonian spiking, we found highly non-trivial effects such as the non-monotonic dependence of the variance of the recurrent input on connection probability (Fig. 5b) and the weak suppression of variability by external stimuli (Fig. 6b). Thus, even though a single population and Poisson spiking dynamics may be too simple to quantitatively model certain phenomena in cortex (such as a stronger suppression of variability [61]), the fact that our MF2 theory explains non-trivial fluctuation effects demonstrates the power of the approach as a proof of principle. Importantly, with the biological extensions discussed above, we believe that our theory may pave the way for a useful modeling framework for variability in cortical circuits at the mesoscopic population level.